Calculate the percent error for your percent by mass of NaClO in the commercial bleach 


Data Table Molarity of Na2S2O3 0.125 M

Initial Buret Reading (Na2S2O3) 1.50 mL

Final Buret Reading (Na2S2O3) 14.35 mL 


1
Expert's answer
2021-01-19T04:29:37-0500

Molarity of Na2S2O3= 0.125M

Volume of Na2S2O3= 14.35ml - 1.50ml= 12.85ml

Molar mass of Na2S2O3= 158gmol-1

Mole= C x V

Mole= 0.125 x 12.85/1000 = 0.0016mol

Mass of Na2S2O3= mole x molar mass

= 0.0016 x 158= 0.25g

Density of the bleach=1.08g/ml

Mass of bleach= density of bleach x Volume

= 1.08 x 12.85=13.88g

Percentage by mass of Na2S2O3 present= 0.25/13.88 x 100% = 1.8%

According to the question, if NaClO is the other substance present in the bleach, then the percentage by mass of NaClO will be = 100 - 1.8 = 98.2%